# Vulnerabilities in AI-generated Image Detection: The Challenge of Adversarial Attacks

### v1

@@ Abstract.
Recent advancements in image synthesis, particularly with the advent of GAN and Diffusion models, have amplified public concerns regarding the dissemination of disinformation. To address such concerns, numerous AI-generated Image (AIGI) Detectors have been proposed and achieved promising performance in identifying fake images. However, there still lacks a systematic understanding of the adversarial robustness of AIGI detectors. In this paper, we examine the vulnerability of state-of-the-art AIGI detectors against adversarial attack under white-box and black-box settings, which has been rarely investigated so far. For the task of AIGI detection, we propose a new attack containing two main parts. First, inspired by the obvious difference between real images and fake images in the frequency domain, we add perturbations under the frequency domain to push the image away from its original frequency distribution. Second, we explore the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous models, e.g. transferring adversarial examples across CNNs and ViTs. This is achieved by introducing a novel post-train Bayesian strategy that turns a single surrogate into a Bayesian one, capable of simulating diverse victim models using one pre-trained surrogate, without the need for re-training. We name our method as frequency-based post-train Bayesian attack, or FPBA. Through FPBA, we show that adversarial attack is truly a real threat to AIGI detectors, because FPBA can deliver successful black-box attacks across models, generators, defense methods, and even evade cross-generator detection, which is a crucial real-world detection scenario. The code will be shared upon acceptance.
AI-generated Image Detection, Adversarial Examples
ccs: Security and privacy Software and application securityccs: Security and privacy Human and societal aspects of security and privacyccs: Computing methodologies Computer vision
1. Introduction
@@
The notable progress in generative models, such as GANs (Goodfellow et al., 2014a) and Diffusion models (Ho et al., 2020), is driving the flourishing development of the image synthesis domain. These generated fake images exhibit realistic-looking, rendering them visually indistinguishable from real images. Moreover, a variety of free and open-source tools facilitates the effortless creation of fake images. However, alongside the benefits, the widespread availability of fake images raises concerns regarding the dissemination of misinformation and fake media news.
@@
Consequently, numerous detectors have been proposed(Durall et al., 2020; Chai et al., 2020; Corvi et al., 2023b) to identify AI-generated images (AIGI). Recent state-of-the-art detectors (Wang et al., 2020; Zhong et al., 2023) rely on Deep Neural Networks (DNNs) to classify, achieving significant accuracy performance across multiple datasets and generative models. However, our investigation reveals that AIGI detectors are vulnerable to adversarial examples, capable of misleading detectors by classifying fake images as real. Despite several works (Carlini and Farid, 2020; Hussain et al., 2021; Neekhara et al., 2021; Hou et al., 2023; Jia et al., 2022a) exploring adversarial attacks for GAN-based face forgery detection, we extend this security issue to encompass broader AIGI detection, which has been largely ignored. Unlike face forgery concentrates solely on manipulating or synthesizing faces, often crafted by GANs, AI-generated images span a broader spectrum, generated by a variety of advanced diffusion models and GANs. These images encompass a richer array of semantic content, thus greatly enhancing their potential for disseminating disinformation.
Figure 1. Illustration of the proposed method.
@@
In this paper, we show that AIGI detectors are vulnerable to adversarial attacks. Considering that many works (Dzanic et al., 2020; Frank et al., 2020) have demonstrated the obvious changes between real and fake images in the frequency domain, we explore the vulnerable region of AIGI detectors in the frequency domain. As illustrated in Fig. 3, the frequency components that different detectors focus on significantly vary from each other. Therefore, we utilize frequency spectrum transformation to uncover diverse substitute models. To further improve the adversarial transferability across heterogeneous models, e.g. transferring adversarial examples across Visual Transformers (ViTs) and Convolutional Neural Networks (CNNs), we propose a post-train Bayesian strategy to conduct a Bayesian treatment on the surrogate model, without the need for re-training the surrogate. In contrast to existing ensemble-based or Bayesian attacks, which involve retraining an ensemble of surrogate models, our post-train Bayesian strategy freezes the pre-trained surrogate and appends tiny extra Bayesian components behind the surrogate, avoiding a heavy memory footprint and speeding up the training process. As a result, we propose the first adversarial attack for general AIGI detection, to add adversarial perturbations in the frequency domain from a post-train Bayesian perspective. We name our method Frequency-based Post-train Bayesian Attack, or FPBA. A high-level illustration of our method and the key differences between our method and the previous method are shown in Fig. 1.
@@
  • To our best knowledge, we are the first to systemically investigate the adversarial vulnerability of state-of-the-art AIGI detectors. We extensively evaluate the adversarial robustness of traditionally trained models, defense models, and cross-generated detection within real-world scenarios. We demonstrate adversarial examples are truly a real threat to AIGI detections, and reveal the phenomenon of gradient masking phenomena in AIGI detectors.
@@
  • We conduct comprehensive experiments on eight spatial-based detectors with CNN-based, Vit-based and CLIP-based architectures, and two frequency-based detectors using AIGI and Deepfake datasets. The extensive results demonstrate our proposed method achieves the highest average attack success rate under white-box and black-box settings, outperforming baseline methods by a big margin.
2. Related Work
@@
AI-Generated Image Detection: The rapid advancement of AI generative technologies, including GANs (Goodfellow et al., 2014a), VAEs (Kingma and Welling, 2014), and Diffusion models (Ho et al., 2020), has brought about substantial risks in the dissemination of disinformation. Initially, early VAEs and GANs models were primarily utilized to generate realistic fake face data, leading detection methods to primarily focus on face forgeries. However, with the emergence of diffusion models, the scope of generative models has expanded to encompass diverse natural scene objects. Consequently, this expansion presents more significant challenges to detection methods, as they must now address a broader range of forgery types and scenarios.
@@
The primary objective of AIGI detection methods is to discern the authenticity of an image by framing it as a binary classification task, distinguishing between real and fake instances. Data-driven-based approaches (Durall et al., 2020; Chai et al., 2020; Wang et al., 2020; Corvi et al., 2023b) have demonstrated promising results on generated data identically distributed. However, their performance significantly degraded when presented with previously unseen generated images. More recently, certain methods (Dzanic et al., 2020; Frank et al., 2020; Chandrasegaran et al., 2021) focus on analyzing forgery traces in the frequency domain and learning relevant forgery features (Corvi et al., 2023a; Liu et al., 2023; Pontorno et al., 2024; Zhang et al., 2019) to enhance the generalization capability of the detection system. In pursuit of these objectives, some methods have been proposed to learn local forgery features (Ju et al., 2023; Lorenz et al., 2023) in the spatial domain. AIGI fingerprint representations (Jeong et al., 2022) are acquired through the learning of noise patterns. Notably, studies such as (Liu et al., 2022; Zhong et al., 2023) have demonstrated the utility of utilizing noise in the frequency domain as a fingerprint for enhancing performance. A sequence of detection methodologies (Tan et al., 2023; Liu et al., 2023; Cozzolino et al., 2023; Chang et al., 2023) has been progressively investigated, relying on the practice of freezing pre-trained models. These approaches employ parameter-efficient fine-tuning strategies while integrating additional classifiers. For instance, UniFD (Ojha et al., 2023) explores pre-trained vision-language models for AIGI detection, showcasing the effectiveness of leveraging such models in the pursuit of robust forgery detection. Furthermore, other researchers have pursued a comprehensive investigation of generalized forgery features from various perspectives, ranging from diffusion reconstruction error (Wang et al., 2023a; Ricker et al., 2024), adversarial teacher-student discrepancy-aware framework (Zhu et al., 2023a) or diffusion noise(Zhang and Xu, 2023).
The existing models consistently demonstrate enhancement in detecting AI-generated images. However, their behaviors under adversarial attacks have not been thoroughly examined. Our work is complementary to existing research by making the first attempt to systemically evaluate their robustness to adversarial attacks and suggesting potential improvements.
@@
Adversarial Attack: Adversarial attacks aim to generate adversarial examples that can fool the target model into predicting wrongly. The vulnerability of DNNs has sparked significant concerns in many safety-critical fields, including face recognition (Jia et al., 2022b), autonomous driving (Zhu et al., 2023b; Xie et al., 2023) and human activity recognition (Diao et al., 2021). Very recently, the adversarial attack against Deepfake detection has drawn attention. Hussain et al. (Hussain et al., 2021) employ gradient sign-based methods (Kurakin et al., 2018; Goodfellow et al., 2014b) to evaluate the adversarial robustness of video Deepfake detectors. Li et al. (Li et al., 2021) generate adversarial fake images along the latent face manifold. Carlini et al. (Carlini and Farid, 2020) and Neekhara et al. (Neekhara et al., 2021) investigate the adversarial threat to Deepfake detection under the black-box setting. Jia et al. (Jia et al., 2022a) propose to generate adversarial perturbations in the frequency domain. Hou et al. (Hou et al., 2023) narrow the feature distribution gap between real and fake images by incorporating natural degradation noise with adversarial perturbations.
Different from Deepfake tampering the real face images, AI-generated image encompasses a broader range of synthetic images with more diverse semantic content, amplifying their potential for dissemination of disinformation. However, the adversarial robustness of AIGI detection has received little attention. In this paper, we propose a universal attack against both AIGI detectors and Deepfake detectors, and evaluate under the white-box and a more practical black-box setting.
3. Methodology
@@
Figure 2. The workflow of FPBA. We add spatial-frequency adversarial perturbations to AI-generated images in a Bayesian manner, so that they are misclassified as real.
3.1. Preliminaries
@@
(1) $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\tilde{\mathbf{x}},\Theta)=\operatorname*{arg\,max}_{\tilde{\mathbf{x}}}L(\tilde{\mathbf{x}},y,\Theta),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon$
where $L$ is the binary cross entropy loss in AI-generated image detection. Adversarial example $\tilde{\mathbf{x}}=\mathbf{x}+\delta$, in which $\delta$ is the adversarial perturbation and $\epsilon$ is the perturbation budget. Eq. 1 can be performed with iterative gradient-based methods, such as PGD (Madry et al., 2017) or I-FGSM (Kurakin et al., 2018):
(2) $\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}+\alpha\cdot\operatorname{sign}(\nabla L\left(\tilde{\mathbf{x}}^{i},y,\Theta\right))$
(3) $=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}(\nabla\,log\,p(y\mid\tilde{\mathbf{x}}^{i},\Theta)).$
3.2. Frequency-based Analysis and Attacks
@@
Many AIGI detection approaches distinguish between real and fake images via subtle artifacts (Wang et al., 2020; Jeong et al., 2022; Zhong et al., 2023). While these subtle clues are invisible in the spatial domain, a series of works (Dzanic et al., 2020; Frank et al., 2020; Chandrasegaran et al., 2021) demonstrate that there are obvious differences between real and fake images in the frequency domain. This inspires us to explore the vulnerable region of AIGI detectors from a frequency perspective. To this end, we first implement discrete cosine transform (DCT) $\mathcal{D}(\cdot)$ to transfer the inputs from the spatial domain to the frequency domain. To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map (Long et al., 2022) to visualize the sensitive components of real and fake images across different models:
(4) $\mathbf{S}_{\Theta}=\frac{\partial J(\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}),y,\Theta)}{\partial\mathcal{D}(\mathbf{x})}$
@@
(5) $\Gamma(\mathbf{x})=\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}+\xi)\odot\mathcal{M}),$
where $\Gamma(\cdot)$ denotes the random spectrum transformation (Long et al., 2022). $\odot$ is the Hadamard product, $\xi$ is a random noise drawn from an isotropic Gaussian $\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right)$, and each element of $\mathcal{M}$ is sampled from a Uniform distribution $\mathcal{U}(1-p,1+p)$. As shown in Fig. 3(a), tuning the spectrum saliency map can cover most of the other models. We hence conduct adversarial attack in the frequency domain via spectrum transformation:
(6) $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\Gamma(\tilde{\mathbf{x}}),\Theta),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon$
3.3. Exploring the Surrogate Posterior Space
Although tuning the spectrum transformation in Eq. 6 can simulate different substitute models with a homogeneous architecture (Long et al., 2022), it shows limited transferability when applied to heterogeneous architectures, e.g. transferring adversarial examples across ViTs and CNNs. This motivates us to consider the frequency-based attack from a Bayesian perspective, i.e. exploring the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous models. Therefore, we redefine Eq. 6 by minimizing the Bayesian posterior predictive distribution:
@@ 3.3. Exploring the Surrogate Posterior Space
(7) $=$ $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}\,\mathbb{E}_{\Theta\sim p(\Theta\mid\mathcal{D})}p\left(y\mid\Gamma(\tilde{\mathbf{x}}),\Theta\right),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon$
@@
Figure 4. The architecture of the appended model. $\sigma$ means the sigmoid layer.
3.3.1. Post-train Bayesian Strategy for Frequency-based Attack
However, it is not straightforward to attack AIGI detectors in such a Bayesian manner due to several factors. First, the Bayesian posterior for DNNs is a high dimensional distribution due to a very large number of parameters of DNNs (Izmailov et al., 2021). Hence computing and sampling the posterior distribution is an intractable problem. Albeit feasible for approximate sampling the posterior via variational inference or Markov Chain Monte Carlo (MCMC), it is computationally slow and expensive in such a high dimensional space. Furthermore, to improve the accuracy and generalization of the AIGI detectors, there is a growing inclination to train AIGI detectors on large-scale datasets (Zhu et al., 2024; He et al., 2021). From the perspective of end-users, it is not desirable to re-train a surrogate model on large-scale datasets for attack.

### v3

@@ Abstract
Recent advancements in image synthesis, particularly with the advent of GAN and Diffusion models, have amplified public concerns regarding the dissemination of disinformation. To address such concerns, numerous AI-generated Image (AIGI) Detectors have been proposed and achieved promising performance in identifying fake images. However, there still lacks a systematic understanding of the adversarial robustness of AIGI detectors. In this paper, we examine the vulnerability of state-of-the-art AIGI detectors against adversarial attack under white-box and black-box settings, which has been rarely investigated so far. To this end, we propose a new method to attack AIGI detectors. First, inspired by the obvious difference between real images and fake images in the frequency domain, we add perturbations under the frequency domain to push the image away from its original frequency distribution. Second, we explore the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous AIGI detectors, e.g. transferring adversarial examples across CNNs and ViTs. This is achieved by introducing a novel post-train Bayesian strategy that turns a single surrogate into a Bayesian one, capable of simulating diverse victim models using one pre-trained surrogate, without the need for re-training. We name our method as Frequency-based Post-train Bayesian Attack, or FPBA. Through FPBA, we show that adversarial attack is truly a real threat to AIGI detectors, because FPBA can deliver successful black-box attacks across models, generators, defense methods, and even evade cross-generator detection, which is a crucial real-world detection scenario. The code will be shared upon acceptance.
Index Terms:
AI-generated Image Detection, Adversarial Examples.
I Introduction
Figure 1: A high-level illustration of our proposed method.
@@ I Introduction
The notable progress in generative models, such as GANs [1] and Diffusion models [2], is driving the flourishing development of the image synthesis domain. These generated fake images exhibit realistic-looking, rendering them visually indistinguishable from real images. Moreover, a variety of free and open-source tools facilitates the effortless creation of fake images. However, alongside the benefits, the widespread availability of fake images raises concerns regarding the dissemination of misinformation and fake media news.
@@
Consequently, numerous detectors have been proposed to identify AI-generated images (AIGI). Recent state-of-the-art detectors [3, 4] rely on Deep Neural Networks (DNNs) to classify, achieving significant accuracy performance across multiple datasets and generative models. However, our investigation reveals that AIGI detectors are vulnerable to adversarial examples, capable of misleading detectors by classifying fake images as real. Despite several works [5, 6, 7, 8, 9] exploring adversarial attacks for GAN-based face forgery detection, we extend this security issue to encompass broader AIGI detection, which has been largely ignored. Unlike face forgery concentrates solely on manipulating or synthesizing faces, often crafted by GANs, AI-generated images span a broader spectrum, generated by a variety of advanced diffusion models and GANs. These images encompass a richer array of semantic content, thus greatly enhancing their potential for disseminating disinformation.
@@
In this paper, we show that AIGI detectors are vulnerable to adversarial attacks. Considering that many works [10, 11] have demonstrated the obvious changes between real and fake images in the frequency domain, we explore the vulnerable region of AIGI detectors in the frequency domain. As illustrated in Fig. 3, the frequency components that different detectors focus on significantly vary from each other. Therefore, we utilize frequency spectrum transformation to uncover diverse substitute models via adding adversarial perturbations in the various frequency transformation domains. Further, Transformer-based detectors have demonstrated outstanding performance in detection AI-generated images[12], but we have observed that there is an obvious gap in adversarial transferability across heterogeneous AIGI detectors e.g. transferring adversarial examples across Convolutional Neural Networks (CNNs) to Visual Transformers (ViTs) (as shown in Table I). To tackle this issue, we propose a post-train Bayesian strategy to conduct a Bayesian treatment on the surrogate model, without the need for re-training the surrogate. In contrast to existing ensemble-based or Bayesian attacks, which involve retraining an ensemble of surrogate models, our post-train Bayesian strategy freezes the pre-trained surrogate and appends tiny extra Bayesian components behind the surrogate, avoiding a heavy memory footprint and speeding up the training process. As a result, we propose a new transferable adversarial attack for general AIGI detection, to add adversarial perturbations in various frequency transformation domains from a post-train Bayesian perspective. We name our method Frequency-based Post-train Bayesian Attack, or FPBA. A high-level illustration of our method and the key differences between our method and the previous method are shown in Fig. 1.
@@
  • We systemically investigate the adversarial robustness of state-of-the-art AIGI detectors, including standard trained models, defense models, and cross-generated detection within real-world scenarios. We demonstrate adversarial examples are truly a real threat to AIGI detections, and reveal the phenomenon of gradient masking phenomena in AIGI detectors.
@@
  • We conduct comprehensive experiments on 13 various AIGI detectors, including heterogeneous model architectures and various detection methods. Extensive results demonstrate that our proposed method achieves the highest average attack success rate under white-box and black-box settings, outperforming baseline methods by a big margin.
II Related Work
@@
AI-Generated Image Detection: The rapid advancement of AI generative technologies, including GANs [1], VAEs [13], and Diffusion models [2], has brought about substantial risks in the dissemination of disinformation. Initially, early VAEs and GANs models were primarily utilized to generate realistic fake face data, leading detection methods [14, 15, 16], to primarily focus on face forgeries. However, with the emergence of diffusion models, the scope of generative models has expanded to encompass diverse natural scene objects. Consequently, this expansion presents more significant challenges to detection methods, as they must now address a broader range of forgery types and scenarios.
@@
The primary objective of AIGI detection methods is to discern the authenticity of an image by framing it as a binary classification task, distinguishing between real and fake instances. Data-driven-based approaches [3, 17] have demonstrated promising results on generated data identically distributed. However, their performance significantly degraded when presented with previously unseen generated images. More recently, certain methods [10, 11] focus on analyzing forgery traces in the frequency domain and learning relevant forgery features [18] to enhance the generalization capability of the detection system. In pursuit of these objectives, some methods have been proposed to learn local forgery features [19, 20] in the spatial domain. AIGI fingerprint representations [21] are acquired through the learning of noise patterns. Notably, studies such as [22, 4] have demonstrated the utility of utilizing noise in the frequency domain as a fingerprint for enhancing performance. A sequence of detection methodologies [23] has been progressively investigated, relying on the practice of freezing pre-trained models. These approaches employ parameter-efficient fine-tuning strategies while integrating additional classifiers. For instance, UniFD [24] explores pre-trained vision-language models for AIGI detection, showcasing the effectiveness of leveraging such models in the pursuit of robust forgery detection. Furthermore, other researchers have pursued a comprehensive investigation of generalized forgery features from various perspectives, ranging from diffusion reconstruction error [25, 26], adversarial teacher-student discrepancy-aware framework [27] or diffusion noise[28]. The existing models consistently demonstrate enhancement in detecting AI-generated images. However, their behaviors under adversarial attacks have not been thoroughly examined. Our work is complementary to existing research by making the attempt to systemically evaluate their robustness to adversarial attacks and suggesting potential improvements.
@@
Adversarial Attack: Adversarial attacks aim to generate adversarial examples that can fool the target model into predicting wrongly. The vulnerability of DNNs has sparked significant concerns in many safety-critical fields, including image classification[29, 30], multimedia communication [31] and human activity recognition [32]. Very recently, the adversarial attack against Deepfake detection has drawn attention. Hussain et al. [6] employ gradient sign-based methods [33, 34] to evaluate the adversarial robustness of video Deepfake detectors. Li et al. [35] generate adversarial fake images along the latent face manifold. Carlini et al. [5] and Neekhara et al. [7] investigate the adversarial threat to Deepfake detection under the black-box setting. Jia et al. [9] propose to generate adversarial perturbations in the frequency domain. Hou et al. [8] narrow the feature distribution gap between real and fake images by incorporating natural degradation noise with adversarial perturbations. FakePolisher [36] and TraceEvader [37] improve the untraceability of deepfakes by removing artificial fingerprints. Different from Deepfake tampering the real face images, AI-generated image encompasses a broader range of synthetic images with more diverse semantic content, amplifying their potential for dissemination of disinformation. However, the adversarial robustness of AIGI detection has received little attention. In this paper, we propose a universal attack against both AIGI detectors and Deepfake detectors, and evaluate under the white-box and a more practical black-box setting.
III Methodology
@@
Figure 2: The workflow of FPBA. We add spatial-frequency adversarial perturbations to AI-generated images in a Bayesian manner, so that they are misclassified as real. DCT and IDCT are discrete cosine transformation and inverse discrete cosine transformation respectively.
III-A Preliminaries
@@
$\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\tilde{\mathbf{x}},\Theta)=\operatorname*{arg\,max}_{\tilde{\mathbf{x}}}L(\tilde{\mathbf{x}},y,\Theta),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon,$ (1)
where $L$ is the binary cross entropy loss in AI-generated image detection. Adversarial example $\tilde{\mathbf{x}}=\mathbf{x}+\delta$, in which $\delta$ is the adversarial perturbation and $\epsilon$ is the perturbation budget. Eq. 1 can be performed with iterative gradient-based methods, such as PGD [38] or I-FGSM [33]:
$\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}+\alpha\cdot\operatorname{sign}(\nabla L\left(\tilde{\mathbf{x}}^{i},y,\Theta\right))$
$=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}(\nabla\,log\,p(y\mid\tilde{\mathbf{x}}^{i},\Theta)).$ (2)
@@
III-B Frequency-based Analysis and Attacks
Many AIGI detection approaches distinguish between real and fake images via subtle artifacts [3, 21, 4]. While these subtle clues are invisible in the spatial domain, a series of works [10, 11] demonstrate that there are obvious differences between real and fake images in the frequency domain. This inspires us to explore the vulnerable region of AIGI detectors from a frequency perspective. To this end, we first implement discrete cosine transform (DCT) $\mathcal{D}(\cdot)$ to transfer the inputs from the spatial domain to the frequency domain. To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map [39] to visualize the sensitive components of real and fake images across different models:
$\mathbf{S}_{\Theta}=\frac{\partial J(\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}),y,\Theta)}{\partial\mathcal{D}(\mathbf{x})},$ (3)
@@ III-B Frequency-based Analysis and Attacks
$\Gamma(\mathbf{x})=\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}+\xi)\odot\mathcal{M}),$ (4)
where $\Gamma(\cdot)$ denotes the random spectrum transformation [39]. $\odot$ is the Hadamard product, $\xi$ is a random noise drawn from an isotropic Gaussian $\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right)$, and each element of $\mathcal{M}$ is sampled from a Uniform distribution $\mathcal{U}(1-p,1+p)$. As shown in Fig. 3(a), tuning the spectrum saliency map can cover most of the other models. We hence conduct adversarial attack in the frequency domain via spectrum transformation:
$\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\Gamma(\tilde{\mathbf{x}}),\Theta),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon.$ (5)
III-C Exploring the Surrogate Posterior Space
Although tuning the spectrum transformation in Eq. 5 can simulate different substitute models with a homogeneous architecture [39], it shows limited transferability when applied to heterogeneous architectures, e.g. transferring adversarial examples across ViTs and CNNs. This motivates us to consider the frequency-based attack from a Bayesian perspective, i.e. exploring the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous models. Therefore, we redefine Eq. 5 by minimizing the Bayesian posterior predictive distribution:
@@ III-C Exploring the Surrogate Posterior Space
$=$ $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}\,\mathbb{E}_{\Theta\sim p(\Theta\mid\mathcal{D})}p\left(y\mid\Gamma(\tilde{\mathbf{x}}),\Theta\right),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon,$ (6)
@@
Figure 4: The architecture of the appended model. $\sigma$ means the sigmoid layer.
III-C1 Post-train Bayesian Strategy
However, it is not straightforward to attack AIGI detectors in such a Bayesian manner due to several factors. First, the Bayesian posterior for DNNs is a high dimensional distribution due to a very large number of parameters of DNNs [40]. Hence computing and sampling the posterior distribution is an intractable problem. Albeit feasible for approximate sampling the posterior via variational inference or Markov Chain Monte Carlo (MCMC), it is computationally slow and expensive in such a high dimensional space. Furthermore, to improve the accuracy and generalization of the AIGI detectors, there is a growing inclination to train AIGI detectors on large-scale datasets [12, 41]. From the perspective of end-users, it is not desirable to re-train a surrogate model on large-scale datasets for attack.